# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@ 3. Few-Shot Learning Models
In this section, we analyze some of the major approaches proposed over the years in the context of FSL. Most existing surveys have mainly focused on meta-learning, transfer learning, and data augmentation techniques. Yet, due to the recent bloom that FSL has seen, we extend this framework to include additional families of methods such as Neural Processes and hybrid approaches that lie at the intersection of multiple disciplines. Our focus is on novel approaches that have surfaced in recent years. Yet, an overview of older but well-established methods is also provided to delineate the background upon which more recent works have been developed. Figure 1 provides an overview of the different types of methods within the FSL paradigm as well as some of their major fields of application.
Figure 1. A taxonomy of FSL methods and their applications.
3.1. Meta-Learning Approaches
@@
Traditionally, meta-learning has been the leading paradigm to deal with FSL. As a result, a wide variety of approaches have been proposed over the years that fall under this category, including metric-based, model-based, optimization-based, and probabilistic methods.
3.1.1. Metric-Based Methods
@@
One of the first FSL models based on metric learning was Siamese neural networks(Koch et al., 2015). Specifically, their objective is to learn how to discriminate between two images, similar to contrastive learning. Under this setup, both support and query set samples are encoded using the same Convolutional Neural Network (CNN) feature extractor, and classification is done based on the $\ell_{1}$ distance between each query set sample’s representation and the representations of the support set samples.
@@
On the other hand, Matching networks(Vinyals et al., 2016) employ different CNN embedding networks for support and query sets. The predicted output is the weighted sum of the support set sample labels, and the weights are calculated using an attention mechanism based on cosine similarity between the query and support set samples’ representations. Full Context Embeddings are also proposed, where support and query set samples are contextualized on the support set embeddings using an LSTM model. Figure 3 shows a high-level overview of Matching Networks’ architecture.
Figure 2. Matching networks architecture (Vinyals et al., 2016).
Figure 3. Prototypical networks architecture(Snell et al., 2017).
@@
Another popular method, Prototypical Networks(Snell et al., 2017), relies on learning a suitable feature extractor whose feature representations of the same class samples can be clustered around a single prototype representation. A single CNN feature extractor is used for both support and query set samples, and each class prototype is the mean of the representations belonging to that class. Next, a simple nearest class prototype algorithm is employed to classify each query sample. Figure 3 shows a high-level overview of Prototypical Networks’ architecture. Recently, Prototypical Networks have been extended in the setting of meta-learning with limited tasks by training them on both real and synthetic tasks. For the creation of synthetic tasks, Conditional Batch Normalization (CNB) is applied to the data of an existing task to create modulated features either using a deterministic objective or via variational inference(Sun et al., 2023).
@@
Similar to Prototypical Networks, Relation Networks(Sung et al., 2018) also use a single feature extractor for both support and query sets and utilize class prototypes. However, in Relation Networks, the nearest class prototype algorithm is replaced by an MLP module that learns with which prototype the query representation is the most similar. In general, this can be seen as a metric learning approach where, instead of using a predefined distance metric (e.g., Euclidean distance in Prototypical Networks), the metric is learned by a neural network. Figure 5 shows a high-level overview of Relation Networks’ architecture.
@@
In (Garcia and Bruna, 2018), FSL is approached by a different perspective where each task is modeled as a partially observed graph. While the feature representations of support and query set samples are still extracted to form the graph, a Graph Neural Network is used to draw inferences, acting as a form of learnable metric. Notably, the proposed method is general enough and includes previous architectures, such as Prototypical Networks and Siamese Networks, as special instances. It can also be applied for semi-supervised and active FSL.
@@
Finally, DeepEMD(Zhang et al., 2020) aims to address the limitations of using global features by leveraging local feature characteristics that provide more discriminative and transferable information across categories. Specifically, instead of comparing representations of whole images, these are broken down into patches, and Earth Mover’s Distance (EMD), which is incorporated as a layer, is used to calculate their similarity. Additionally, a cross-reference mechanism is used to discard patches that introduce noise due to cluttered backgrounds and large intra-class appearance variations.
@@
One of the first works to address the limitations of using task-independent feature extractors was TADAM(Oreshkin et al., 2018), which introduced the use of an extractor contextualized on each task’s data. Contrary to Relation Networks’ representation contextualization that is done as a post-processing step, in TADAM, the feature extractor is directly modulated via learnable task-specific shift and scale vectors for each convolutional layer, and a learnable scale parameter is introduced in the model’s loss function. Finally, metric scaling is proposed to bridge the gap between using different metrics.
FEAT(Ye et al., 2020) also leverages task-specific representations by contextualizing each support set sample to the rest of the set’s samples using a self-attention mechanism, while a contrastive objective is incorporated into the loss function to enhance discrimination between different classes. A transformer-based architecture has also been used in Sparse Spatial Transformers(Chen et al., 2023b) (SSFormers). In particular, SSFormers aim to improve DeepEMD by contextualizing an image’s local features to the rest of the image, as well as patches belonging to other support set images, by using sparse attention maps of the extracted task-relevant patches.
Recently, various task-specific adaptations of existing task-invariant methods have also been proposed. For instance, finetuning of Prototypical Networks’ backbone feature extractor for each task has been proposed, based on the observation that learned feature representation may fail to generalize to novel tasks due to possible distributional shifts(Hu et al., 2022a). Additionally, in (Kim et al., 2023), a model that can adapt to different dense prediction tasks by matching similar image patches and using the weighted mean of the output patches as the prediction is proposed. This model can be seen as a generalization of Matching Networks where matching is performed on the latent representations of the labels derived from similar image patches.
Finally, in (Hu et al., 2022b), the problem of few-shot semantic segmentation is addressed by learning a feature extractor that addresses different sources of heterogeneity. In particular, it incorporates a cross-sample attention module, the fusion of information between different background patches, and a masked image segmentation scheme to reinforce the capacity of contextual inference. Contrary to previously mentioned approaches, this method focuses on learning more separable feature representations by enhancing intra-class compactness and leveraging their distances during training.
3.1.2. Model-Based Methods
@@
Some of the earliest works in this line of research have been primarily based on the fact that the hidden states of Recurrent Neural Networks (RNNs) can be utilized to model learning algorithms. For instance, in (Duan et al., 2016), learning a reinforcement learning (RL) algorithm capable of rapidly adapting to novel tasks is also formulated as an RL problem. Based on the concept that in meta-learning, knowledge acquisition takes place in two different time scales, a GRU network’s internal state is used as a fast learner to adapt to a new task with few samples, while Trust Region Policy Optimization is used to optimize the slow learner’s policy across tasks. Consequently, the model can generalize within different tasks like multi-armed bandits, tabular Markov Decision Processes (MDPs), and visual navigation (partially observable MDPs). On the other hand, in (Wang et al., 2016), a similar approach based on RL is applied for structured task distributions. The proposed deep meta-RL approach trains an RNN to act as a learning algorithm that can be applied to different MDPs from the same distribution. Using reward information from various tasks, the RNN can quickly adjust to new tasks during testing without finetuning due to its internal state and ability to model different learning algorithms.
On a more general approach, SNAIL(Mishra et al., 2018) can be applied to different types of problems without any assumptions for the task or the optimization procedure. It uses a generic architecture based on three building blocks: dense blocks (temporal 1D causal convolutions), TCBlocks (dense blocks with increasing dilation rates), and a self-attention block. Interestingly, using temporal convolutions interleaved with attention enables selecting relevant information over a larger context window. Overall, this approach allows for learning an optimal strategy without any built-in assumptions, as in gradient-based meta-learning.
Finally, some recent approaches have introduced architecture-based models suitable for or inspired by different domains, such as a meta-learning model for forecasting dynamical systems that generalizes across different time series sharing a common structure(Wang et al., 2022a). This is achieved by extracting time-invariant features of each time series dynamics using a CNN-based encoder and then adapting to the dynamics of each specific task via adaptive modules. On the other hand, inspired by the Lottery Ticket Hypothesis(Frankle and Carbin, 2018), a gradient-based meta-learning method is proposed in (Chijiwa et al., 2022) to meta-learn a sparse neural network structure suitable for FSL. Specifically, the method aims to learn a parameter mask to derive a sparse subnetwork that demonstrates strong meta-generalization capabilities. This method can be seen as a generalization for architecture-based modeling from hand-crafted architectures to learnable ones.
@@
External memory modules can store different types of content that facilitate rapid adaptation, such as model weights or sample representations. Meta Networks(Munkhdalai and Yu, 2017) is a model that uses external memory modules to store fast weights for rapid adaptation across tasks. The model consists of a meta-learner (learning across tasks) and a base-learner (task-specific learning) that are both augmented using additional parameters (fast weights) generated on the fly. These fast weights are used to parameterize each module to rapidly adapt to specific tasks and specific task samples. An external memory module is also used to store these fast weights and retrieve their optimal combination based on the task’s input representations.
On the other hand, most memory-based models typically rely on external memories to store refined sample representations that enable fast adaptation to classes with few samples. In (Santoro et al., 2016), the stored values are retrieved using soft attention and writing to memory is controlled by a content-based memory writer that writes memories to either the least used or the most recently used memory location. Figure 5 shows a high-level overview of the model’s architecture. Additionally, in (Kaiser et al., 2017), an independent memory module is proposed that allows neural networks to achieve one-shot and life-long learning. Specifically, it uses a memory that stores class labels, and the main network’s layer activations are used to query the memory based on their cosine similarity with the memory keys. An approximate nearest neighbors algorithm is used to scale the memory, and the memory can be combined with various network types, leading to improved classification of rare samples.
Figure 4. Relation networks architecture(Sung et al., 2018).
Figure 5. Memory-Augmented Neural Networks architecture(Santoro et al., 2016).
Recently, memory-augmented networks have also been utilized in different applications where rare classes are common. In (Gkanatsios et al., 2023), a memory-augmented 3D object scene segmentation network is proposed. The model consists of a retriever that finds the most relevant representations in the memory and a modulator where the retrieved memories are encoded and enhanced by scene-agnostic queries, while attention modules are also used for contextualization of the query images. Finally, in (Schimunek et al., 2023), a memory-enhanced model for few-shot drug discovery is proposed. Specifically, it uses a context module based on Modern Hopfield Networks (MHNs) to extract representations of the support set and the query set conditioned on a large external context set of molecules via an attention mechanism. Attention is also used to contextualize the support and query sets, and overall robustness is enhanced by reinforcing the covariance structure and reducing the noise.
3.1.3. Optimization-Based Methods
@@
Based on the principle that learned features are better than hand-crafted ones, in this setting, using hand-crafted update rules such as Stochastic Gradient Descent (SGD) is replaced by learned rules. This is achieved by casting the design of the rule as a learning problem. One of the first approaches towards that direction has used LSTMs to output the updates of the gradient descent step used within each task(Andrychowicz et al., 2016). To control all steps of the optimization procedure, an objective dependent on the entire trajectory of learned optimization is used. LSTMs are suitable for this since they maintain past information in their internal state, allowing for previous updates to inform subsequent ones. Meta-Learner LSTM(Ravi and Larochelle, 2016) extends this framework to the FSL setting by learning both the base learner’s initialization and update strategy using only a small, predefined number of steps. Specifically, each LSTM cell state represents the parameters of the learner, and the cell state variations correspond to model updates. However, this method does not scale well with the base learner’s size, and both methods fail to capture any dependencies between different model parameters. Finally, finding an optimal optimization algorithm has also been cast as an RL problem(Li and Malik, 2017). Specifically, learning the optimal way to obtain update steps can be seen as selecting an optimal update policy, which is achieved via guided policy search. A specialized cost function is also used that encourages learning update rules that converge fast.
@@
To address hand-designed optimization algorithms’ slow convergence, some optimization-based algorithms focus on learning initial parameters that converge to optimal parameters with only a few gradient descent iterations. Model-Agnostic Meta-Learning(Finn et al., 2017) (MAML) is the first work that aims to learn a model parameter initialization that allows for rapid adaptation to each given task. In particular, meta-learning is cast as a bi-level optimization problem with two learning loops in two different levels. In the inner loop, the base learner is responsible for adapting the initial parameters in the task’s support set and evaluating them on the task’s query set. In the outer loop, the meta-learner optimizes the initial model parameters using the query set errors across training tasks. As a result, the shared information across tasks is encapsulated in the model’s learned initial parameters, which act as a prior that obviates the need for many optimization steps within each task.
@@
Following MAML, many different extensions and variants have been proposed, aiming to address its limitations. For instance, MAML++(Antoniou et al., 2019) is such an extension that introduces various novel features that aim to enhance MAML’s performance and ensure its training stability, such as an outer-loop loss that is dependent on all of the inner loop steps’ losses, improved batch normalization for meta-learning, and learned inner loop learning rates. The difficulty of choosing optimal learning rates for MAML due to its training instability is also addressed in (Behl et al., 2019), where the authors propose to learn inner- and outer-loop learning rates using hypergradient descent, which adaptively updates the learning rates through gradient descent. Meanwhile, in (Rajeswaran et al., 2019), an implicit differentiation approach that removes the need for differentiating through the inner loop’s optimization path is proposed to address the high computational cost and memory burden imposed by using higher-order derivatives in MAML. Finally, REPTILE(Nichol et al., 2018) introduces a first-order approximation of MAML that obviates the need to backpropagate through inner loop gradients. Consequently, this mitigates the need to calculate second-order derivatives and significantly reduces computational complexity.
@@
Multiple approaches have also focused on improving MAML’s performance and extending its capabilities. In (Rusu et al., 2019), the proposed method decouples optimization-based meta-learning from the high-dimensional space of the model parameters by embedding the model parameters to a lower dimensional latent space. The produced decoded initial parameters are specific for each task and conditioned in its support set, and MAML uses the latent variables as input. Lastly, by formulating tasks as discriminating between in-distribution and out-of-distribution samples for a specific class, MAML has also been extended in detecting out-of-distribution samples(Jeong and Kim, 2020).
@@
The recent advances in theoretical aspects of meta-learning have also led to the development of various theory-inspired MAML adaptations. In (Abbas et al., 2022), to avoid the sharp local minima of MAML’s loss function, sharpness-aware minimization, a method that simultaneously minimizes the loss value and the loss sharpness, is integrated into MAML. Additionally, in (Wang et al., 2023c), to improve both meta-generalization and generalization performance within each task after adaptation, standard regularization is applied in the outer loop, and inverted regularization is applied in the inner loop, increasing adaptation difficulty within each task and forcing the model to learn better-generalized hypotheses at the meta-level.
@@
Learning the update rule in FSL can lead to update rules with non-converging behavior and lacking the inductive bias of gradient descent. To address these limitations, a branch of optimization-based approaches focuses on learning how to control the updates of existing optimizers through a learnable component that modulates the loss gradient (gradient preconditioning). One of the first approaches in that direction extends MAML by additionally meta-learning the learning rates and learning directions for each parameter of the base learner(Li et al., 2017). This allows for more expressive learning but doubles the amount of meta-learned parameters, resulting in higher computational complexity. Gradient preconditioning can also take the form of learning which parameters are shared across tasks and which ones are task-specific by using a learnable mask and transforming the model parameters using meta-learned transformation parameters(Lee and Choi, 2018). This can be seen as extending MAML to learn a parameter subspace for task-specific adaptation using a meta-learned distance metric. In (Park and Oliva, 2019), MAML is extended to meta-learn curvatures. These curvatures are used to transform gradients before passing through the optimizers, thus constituting a form of preconditioning. Preconditioning can also be induced implicitly by introducing nonlinear layers interleaved with the base learner’s layers(Flennerhag et al., 2020). In this case, the occurring preconditioning matrix is meta-learned without backpropagation over the base learner’s parameters, leading to increased efficiency and better scaling. Finally, using a preconditioner in MAML’s inner loop to enable more efficient learning by reshaping the geometry of the parameter space has also been proposed(Kang et al., 2023a). The preconditioner is task-specific and formulated as a Riemannian metric, equivalent to natural gradient descent, and it alleviates meta-overfitting by relying more on individual task characteristics.
3.1.4. Probabilistic Extensions

### v3

@@ 3 Few-Shot Learning Models
In this section, we analyze some of the major approaches proposed over the years in the context of FSL. Most existing surveys have mainly focused on meta-learning, transfer learning, and data augmentation techniques. Yet, due to the recent bloom that FSL has seen, we extend this framework to include additional families of methods such as neural processes under the meta-learning methodologies, and the emergent in-context learning which is commonly found in the recent large language model (LLM) literature. Our focus is on novel approaches that have surfaced in recent years. Yet, an overview of older but well-established methods is also provided to delineate the background upon which more recent works have been developed. Figure 1 provides an overview of the different types of methods within the FSL paradigm as well as some of their major fields of application.
Figure 1: A taxonomy of FSL methods and their applications. This includes main methods focused on supervised FSL settings (Section 3), hybrid methods for FSL beyond the supervised setting (Section 4), and FSL fields of application (Section 5)
3.1 Meta-Learning Approaches
@@
Traditionally, meta-learning has been the leading paradigm to deal with FSL. As a result, a wide variety of approaches have been proposed over the years that fall under this category, including metric-based, model-based, optimization-based, and probabilistic methods, as well as neural processes.
3.1.1 Metric-Based Methods
@@
One of the first FSL models based on metric learning was Siamese Neural Networks [109]. Specifically, their objective is to learn how to discriminate between two images, similar to contrastive learning. Under this setup, both support and query set samples are encoded using the same Convolutional Neural Network (CNN) feature extractor, and classification is done based on the $\ell_{1}$ distance between each query set sample’s representation and the representations of the support set samples.
@@
On the other hand, Matching Networks [205] employ different CNN embedding networks for support and query sets. The predicted output is the weighted sum of the support set sample labels, and the weights are calculated using an attention mechanism based on cosine similarity between the query and support set samples’ representations. Full Context Embeddings are also proposed, where support and query set samples are contextualized on the support set embeddings using an LSTM model. Figure 3 shows a high-level overview of Matching Networks’ architecture.
Figure 2: Matching networks architecture [205]. Predictions are made using an attention mechanism to compare the representations of the support and query set samples
Figure 3: Prototypical networks architecture [187]. A mean representation is calculated for each class, and the query sample is assigned to the class of the nearest centroid
@@
Another popular method, Prototypical Networks [187], relies on learning a suitable feature extractor whose feature representations of the same class samples can be clustered around a single prototype representation. A single CNN feature extractor is used for both support and query set samples, and each class prototype is the mean of the representations belonging to that class. Next, a simple nearest class prototype algorithm is employed to classify each query sample. Figure 3 shows a high-level overview of Prototypical Networks’ architecture. Recently, Prototypical Networks have been extended in the setting of meta-learning with limited tasks by training them on both real and synthetic tasks. For the creation of synthetic tasks, Conditional Batch Normalization (CNB) is applied to the data of an existing task to create modulated features either using a deterministic objective or via variational inference [192].
@@
Similar to Prototypical Networks, Relation Networks [195] also use a single feature extractor for both support and query sets and utilize class prototypes. However, in Relation Networks, the nearest class prototype algorithm is replaced by an MLP module that learns with which prototype the query representation is the most similar. In general, this can be seen as a metric learning approach where, instead of using a predefined distance metric (e.g., Euclidean distance in Prototypical Networks), the metric is learned by a neural network. Figure 5 shows a high-level overview of Relation Networks’ architecture.
@@
In [61], FSL is approached by a different perspective where each task is modeled as a partially observed graph. While the feature representations of support and query set samples are still extracted to form the graph, a Graph Neural Network is used to draw inferences, acting as a form of learnable metric. Notably, the proposed method is general enough and includes previous architectures, such as Prototypical Networks and Siamese Networks, as special instances. It can also be applied for semi-supervised and active FSL.
@@
Finally, DeepEMD [247] aims to address the limitations of using global features by leveraging local feature characteristics that provide more discriminative and transferable information across categories. Specifically, instead of comparing representations of whole images, these are broken down into patches, and Earth Mover’s Distance (EMD), which is incorporated as a layer, is used to calculate their similarity. Additionally, a cross-reference mechanism is used to discard patches that introduce noise due to cluttered backgrounds and large intra-class appearance variations.
@@
One of the first works to address the limitations of using task-independent feature extractors was TADAM [156], which introduced the use of an extractor contextualized on each task’s data. Contrary to Relation Networks’ representation contextualization that is done as a post-processing step, in TADAM, the feature extractor is directly modulated via learnable task-specific shift and scale vectors for each convolutional layer, and a learnable scale parameter is introduced in the model’s loss function. Finally, metric scaling is proposed to bridge the gap between using different metrics.
FEAT [239] also leverages task-specific representations by contextualizing each support set sample to the rest of the set’s samples using a self-attention mechanism, while a contrastive objective is incorporated into the loss function to enhance discrimination between different classes. A transformer-based architecture has also been used in Sparse Spatial Transformers [25] (SSFormers). In particular, SSFormers aim to improve DeepEMD by contextualizing an image’s local features to the rest of the image, as well as patches belonging to other support set images, by using sparse attention maps of the extracted task-relevant patches.
Recently, various task-specific adaptations of existing task-invariant methods have also been proposed. For instance, finetuning of Prototypical Networks’ backbone feature extractor for each task has been proposed, based on the observation that learned feature representation may fail to generalize to novel tasks due to possible distributional shifts [81]. Additionally, in [107], a model that can adapt to different dense prediction tasks by matching similar image patches and using the weighted mean of the output patches as the prediction is proposed. This model can be seen as a generalization of Matching Networks where matching is performed on the latent representations of the labels derived from similar image patches.
Finally, in [82], the problem of few-shot semantic segmentation is addressed by learning a feature extractor that addresses different sources of heterogeneity. In particular, it incorporates a cross-sample attention module, the fusion of information between different background patches, and a masked image segmentation scheme to reinforce the capacity of contextual inference. Contrary to previously mentioned approaches, this method focuses on learning more separable feature representations by enhancing intra-class compactness and leveraging their distances during training.
Figure 4: Relation networks architecture [195]. Support and query set representations are concatenated and classified based on their similarity following a metric learning approach
Figure 5: Memory-Augmented Neural Networks architecture [177]. An external memory module is utilized to store and retrieve refined sample representations within each task
3.1.2 Model-Based Methods
@@
Some of the earliest works in this line of research have been primarily based on the fact that the hidden states of Recurrent Neural Networks (RNNs) can be utilized to model learning algorithms. For instance, in [44], learning a reinforcement learning (RL) algorithm capable of rapidly adapting to novel tasks is also formulated as an RL problem. Based on the concept that in meta-learning, knowledge acquisition takes place in two different time scales, a GRU network’s internal state is used as a fast learner to adapt to a new task with few samples, while Trust Region Policy Optimization is used to optimize the slow learner’s policy across tasks. Consequently, the model can generalize within different tasks like multi-armed bandits, tabular Markov Decision Processes (MDPs), and visual navigation (partially observable MDPs). On the other hand, in [213], a similar approach based on RL is applied for structured task distributions. The proposed deep meta-RL approach trains an RNN to act as a learning algorithm that can be applied to different MDPs from the same distribution. Using reward information from various tasks, the RNN can quickly adjust to new tasks during testing without finetuning due to its internal state and ability to model different learning algorithms.
On a more general approach, SNAIL [146] can be applied to different types of problems without any assumptions for the task or the optimization procedure. It uses a generic architecture based on three building blocks: dense blocks (temporal 1D causal convolutions), TCBlocks (dense blocks with increasing dilation rates), and a self-attention block. Interestingly, using temporal convolutions interleaved with attention enables selecting relevant information over a larger context window. Overall, this approach allows for learning an optimal strategy without any built-in assumptions, as in gradient-based meta-learning.
Finally, some recent approaches have introduced architecture-based models suitable for or inspired by different domains, such as a meta-learning model for forecasting dynamical systems that generalizes across different time series sharing a common structure [215]. This is achieved by extracting time-invariant features of each time series dynamics using a CNN-based encoder and then adapting to the dynamics of each specific task via adaptive modules. On the other hand, inspired by the Lottery Ticket Hypothesis [59], a gradient-based meta-learning method is proposed in [29] to meta-learn a sparse neural network structure suitable for FSL. Specifically, the method aims to learn a parameter mask to derive a sparse subnetwork that demonstrates strong meta-generalization capabilities. This method can be seen as a generalization for architecture-based modeling from hand-crafted architectures to learnable ones.
@@
External memory modules can store different types of content that facilitate rapid adaptation, such as model weights or sample representations. Meta Networks [147] is a model that uses external memory modules to store fast weights for rapid adaptation across tasks. The model consists of a meta-learner (learning across tasks) and a base-learner (task-specific learning) that are both augmented using additional parameters (fast weights) generated on the fly. These fast weights are used to parameterize each module to rapidly adapt to specific tasks and specific task samples. An external memory module is also used to store these fast weights and retrieve their optimal combination based on the task’s input representations.
On the other hand, most memory-based models typically rely on external memories to store refined sample representations that enable fast adaptation to classes with few samples. In [177], the stored values are retrieved using soft attention and writing to memory is controlled by a content-based memory writer that writes memories to either the least used or the most recently used memory location. Figure 5 shows a high-level overview of the model’s architecture. Additionally, in [97], an independent memory module is proposed that allows neural networks to achieve one-shot and life-long learning. Specifically, it uses a memory that stores class labels, and the main network’s layer activations are used to query the memory based on their cosine similarity with the memory keys. An approximate nearest neighbors algorithm is used to scale the memory, and the memory can be combined with various network types, leading to improved classification of rare samples.
Recently, memory-augmented networks have also been utilized in different applications where rare classes are common. In [67], a memory-augmented 3D object scene segmentation network is proposed. The model consists of a retriever that finds the most relevant representations in the memory and a modulator where the retrieved memories are encoded and enhanced by scene-agnostic queries, while attention modules are also used for contextualization of the query images. Finally, in [179], a memory-enhanced model for few-shot drug discovery is proposed. Specifically, it uses a context module based on Modern Hopfield Networks (MHNs) to extract representations of the support set and the query set conditioned on a large external context set of molecules via an attention mechanism. Attention is also used to contextualize the support and query sets, and overall robustness is enhanced by reinforcing the covariance structure and reducing the noise.
3.1.3 Optimization-Based Methods
@@
Based on the principle that learned features are better than hand-crafted ones, in this setting, using hand-crafted update rules such as Stochastic Gradient Descent (SGD) is replaced by learned rules. This is achieved by casting the design of the rule as a learning problem. One of the first approaches towards that direction has used LSTMs to output the updates of the gradient descent step used within each task [7]. To control all steps of the optimization procedure, an objective dependent on the entire trajectory of learned optimization is used. LSTMs are suitable for this since they maintain past information in their internal state, allowing for previous updates to inform subsequent ones. Meta-Learner LSTM [172] extends this framework to the FSL setting by learning both the base learner’s initialization and update strategy using only a small, predefined number of steps. Specifically, each LSTM cell state represents the parameters of the learner, and the cell state variations correspond to model updates. However, this method does not scale well with the base learner’s size, and both methods fail to capture any dependencies between different model parameters. Finally, finding an optimal optimization algorithm has also been cast as an RL problem [123]. Specifically, learning the optimal way to obtain update steps can be seen as selecting an optimal update policy, which is achieved via guided policy search. A specialized cost function is also used that encourages learning update rules that converge fast.
@@
To address hand-designed optimization algorithms’ slow convergence, some optimization-based algorithms focus on learning initial parameters that converge to optimal parameters with only a few gradient descent iterations. Model-Agnostic Meta-Learning [54] (MAML) is the first work that aims to learn a model parameter initialization that allows for rapid adaptation to each given task. In particular, meta-learning is cast as a bi-level optimization problem with two learning loops in two different levels. In the inner loop, the base learner is responsible for adapting the initial parameters in the task’s support set and evaluating them on the task’s query set. In the outer loop, the meta-learner optimizes the initial model parameters using the query set errors across training tasks. As a result, the shared information across tasks is encapsulated in the model’s learned initial parameters, which act as a prior that obviates the need for many optimization steps within each task.
@@
Following MAML, many different extensions and variants have been proposed, aiming to address its limitations. For instance, MAML++ [8] is such an extension that introduces various novel features that aim to enhance MAML’s performance and ensure its training stability, such as an outer-loop loss that is dependent on all of the inner loop steps’ losses, improved batch normalization for meta-learning, and learned inner loop learning rates. The difficulty of choosing optimal learning rates for MAML due to its training instability is also addressed in [15], where the authors propose to learn inner- and outer-loop learning rates using hypergradient descent, which adaptively updates the learning rates through gradient descent. Meanwhile, in [169], an implicit differentiation approach that removes the need for differentiating through the inner loop’s optimization path is proposed to address the high computational cost and memory burden imposed by using higher-order derivatives in MAML. Finally, REPTILE [154] introduces a first-order approximation of MAML that obviates the need to backpropagate through inner loop gradients. Consequently, this mitigates the need to calculate second-order derivatives and significantly reduces computational complexity.
@@
Multiple approaches have also focused on improving MAML’s performance and extending its capabilities. In [176], the proposed method decouples optimization-based meta-learning from the high-dimensional space of the model parameters by embedding the model parameters to a lower dimensional latent space. The produced decoded initial parameters are specific for each task and conditioned in its support set, and MAML uses the latent variables as input. Lastly, by formulating tasks as discriminating between in-distribution and out-of-distribution samples for a specific class, MAML has also been extended in detecting out-of-distribution samples [90].
@@
The recent advances in theoretical aspects of meta-learning have also led to the development of various theory-inspired MAML adaptations. In [1], to avoid the sharp local minima of MAML’s loss function, sharpness-aware minimization, a method that simultaneously minimizes the loss value and the loss sharpness, is integrated into MAML. Additionally, in [214], to improve both meta-generalization and generalization performance within each task after adaptation, standard regularization is applied in the outer loop, and inverted regularization is applied in the inner loop, increasing adaptation difficulty within each task and forcing the model to learn better-generalized hypotheses at the meta-level.
@@
Learning the update rule in FSL can lead to update rules with non-converging behavior and lacking the inductive bias of gradient descent. To address these limitations, a branch of optimization-based approaches focuses on learning how to control the updates of existing optimizers through a learnable component that modulates the loss gradient (gradient preconditioning). One of the first approaches in that direction extends MAML by additionally meta-learning the learning rates and learning directions for each parameter of the base learner [127]. This allows for more expressive learning but doubles the amount of meta-learned parameters, resulting in higher computational complexity. Gradient preconditioning can also take the form of learning which parameters are shared across tasks and which ones are task-specific by using a learnable mask and transforming the model parameters using meta-learned transformation parameters [122]. This can be seen as extending MAML to learn a parameter subspace for task-specific adaptation using a meta-learned distance metric. In [159], MAML is extended to meta-learn curvatures. These curvatures are used to transform gradients before passing through the optimizers, thus constituting a form of preconditioning. Preconditioning can also be induced implicitly by introducing nonlinear layers interleaved with the base learner’s layers [57]. In this case, the occurring preconditioning matrix is meta-learned without backpropagation over the base learner’s parameters, leading to increased efficiency and better scaling. Finally, using a preconditioner in MAML’s inner loop to enable more efficient learning by reshaping the geometry of the parameter space has also been proposed [100]. The preconditioner is task-specific and formulated as a Riemannian metric, equivalent to natural gradient descent, and it alleviates meta-overfitting by relying more on individual task characteristics.
3.1.4 Probabilistic Extensions